OpenStudy (anonymous):

Find five numbers such that Ix*yI=3 (where x and Y are elements of natural numbers sets.

myininaya (myininaya):

x=1,y=3 five numbers?

myininaya (myininaya):

five natural numbers*

OpenStudy (anonymous):

yeaaaa... i got those

myininaya (myininaya):

thats impossible to find 5 natural numbers

OpenStudy (anonymous):

:(

myininaya (myininaya):

that satisfy that

myininaya (myininaya):

* means multiply right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for any natural number n>=1 define An as An={(x,y) e NxN: x*y=n. explain why is symmetric find 5 different numbers such that IAnI=3

myininaya (myininaya):

interesting maybe my interpretation of this is wrong let me see if i can find james

OpenStudy (anonymous):

ok

myininaya (myininaya):

coolsector do you have any input?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no i think the same

OpenStudy (anonymous):

guess we dont really understand "(x,y) e NxN"

myininaya (myininaya):

(x,y) is an element of the cartesian product N X N

myininaya (myininaya):

but i still don't understand why they are asking for 5 natural numbers that satisfy |x*y|=3

myininaya (myininaya):

i don't see how this * could mean multiplication with 5 natural numbers out there to satisfy |x*y|=3

OpenStudy (anonymous):

yes it is not possible : xy = 3 or xy = -3 x = y/3 x = -3/y about x = -3/y there is no need to talk it is negative x = y / 3 only 1,3

OpenStudy (anonymous):

ok,,, i have x=1,x=3 y=1, y=3 i need one more

OpenStudy (anonymous):

oh no my answer is wrong,, because x and y cannot be equal the will be anti symmetric

OpenStudy (anonymous):

:9

OpenStudy (anonymous):

what happened?

OpenStudy (jamesj):

Well actually there's quite a few ordered pairs. All the integral solutions of |xy| = 3 are (x,y) = (1,3) or (1,-3) or (-1,3) or (-1,-3) (3,1) or (-3,1) or (3,-1) or (-3,-1)

myininaya (myininaya):

x and y are natural

OpenStudy (jamesj):

ahh, sorry. I thought they were integers. Then no. There's only solution out of those eight: (x,y) = (1,3) or (3,1)

OpenStudy (anonymous):

:(

myininaya (myininaya):

* so this couldn't be any other operation

myininaya (myininaya):

you are like totally for sure it is meant to be multiplication?

OpenStudy (anonymous):

ok, and i think x and y cannot be the same because if (x,y)e N and (y,x)e N.. is called anti symmetric if x=y

OpenStudy (anonymous):

yes ill take a picture wait

myininaya (myininaya):

ok lol

myininaya (myininaya):

i believe you

OpenStudy (anonymous):

suppose this is a take home test,, but is like burn at home test

OpenStudy (anonymous):

I wish my handwriting was as good as that :-(

myininaya (myininaya):

oh wait a minute

OpenStudy (jamesj):

No, no.

OpenStudy (jamesj):

The exam is asking a completely different question.

myininaya (myininaya):

yep

OpenStudy (anonymous):

:o

OpenStudy (jamesj):

|An| means the cardinality of the set An; i.e., the number of members of An.

myininaya (myininaya):

we want to find n such that the cardinality of A_n is 3

OpenStudy (anonymous):

what?

myininaya (myininaya):

A_4={(1,4),(2,2),(4,1)}

myininaya (myininaya):

thats one n=4

OpenStudy (jamesj):

exactly. But knowing now this is an exam I think it's unethical for you to ask and unethical for us to answer the question, but we can explain the question. as myininaya is doing.

myininaya (myininaya):

now we need 4 more

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

ok james :(

OpenStudy (anonymous):

he didnt even explain,,,

myininaya (myininaya):

so i give you one example you can find some others right?

OpenStudy (anonymous):

yeaa.. let me see...

OpenStudy (anonymous):

:)... i know now what i can do

OpenStudy (anonymous):

thanks u guys

myininaya (myininaya):

lol maybe i shouldn't have said i believe you

OpenStudy (anonymous):

jajajaj,, lol

OpenStudy (anonymous):

i didnt believe myself ;)

myininaya (myininaya):

ok good luck on your test

OpenStudy (anonymous):

thanks